# Sprecher Networks: A Parameter-Efficient Kolmogorov-Arnold Architecture

### v1

@@ 3.2 Optional enhancements
  • Residual connections (Cyclic): When enabled, residual connections are implemented using a dimension-adaptive modulo-based cyclic assignment that maintains architectural coherence. Like convolutional layers that achieve parameter efficiency through local connectivity patterns, cyclic residuals use periodic patterns to reduce parameters from $O(d_{\mathrm{in}}\times d_{\mathrm{out}})$ to $O(\max(d_{\mathrm{in}},d_{\mathrm{out}}))$. This method adapts based on dimensional relationships between adjacent layers:
@@
  • Output affine head: The network may include a learnable affine head applied to the final network output: $f_{\text{final}}(\mathbf{x})=s_{\text{out}}\,f(\mathbf{x})+b_{\text{out}}$, where $s_{\text{out}},b_{\text{out}}\in\mathbb{R}$ (2 parameters) are shared across all output dimensions. In our experiments we initialize $b_{\text{out}}$ to the mean of the training targets (averaged over all outputs when $m>1$) and set $s_{\text{out}}=0.1$.
@@
3.2.1 Toy-$4\to 5$ synthetic regression task
We additionally consider a higher-dimensional vector-valued regression benchmark, Toy-$4\to 5$, with inputs $\mathbf{x}=(x_{1},x_{2},x_{3},x_{4})\in[0,1]^{4}$ and outputs $f(\mathbf{x})\in\mathbb{R}^{5}$ defined by
$f_{1}(\mathbf{x})$ $=\sin(2\pi x_{1})\cos(\pi x_{2}),$
$f_{2}(\mathbf{x})$ $=\exp\!\bigl(-2(x_{1}^{2}+x_{2}^{2})\bigr),$
$f_{3}(\mathbf{x})$ $=x_{3}^{3}-x_{4}^{2}+\tfrac{1}{2}\sin(5x_{3}),$
$f_{4}(\mathbf{x})$ $=\operatorname{sigmoid}\!\bigl(3(x_{1}+x_{2}-x_{3}+x_{4})\bigr),$
$f_{5}(\mathbf{x})$ $=\tfrac{1}{2}\sin(4\pi x_{1}x_{4})+\tfrac{1}{2}\cos(3\pi x_{2}x_{3}),$
where $\operatorname{sigmoid}(t)=\frac{1}{1+e^{-t}}$. Unless otherwise stated, training inputs are sampled i.i.d. uniformly from $0,1]^{4}$.
Table 1: Cyclic residuals vs. linear residual projections on synthetic regression tasks. We report the best MSE on a fixed 1,024-point training set (measured in evaluation mode) achieved within the stated number of epochs. Toy-2D-Complex refers to the scalar 2D regression benchmark defined above; Toy-2D (vector) refers to the vector-valued 2D benchmark in Fig. 9; Toy-$4\to 5$ refers to the synthetic regression task defined in Section 3.2.1. “Linear” uses a learned scalar when $d_{\mathrm{in}}=d_{\mathrm{out}}$ and a learned matrix $W\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ otherwise; “cyclic” uses the pooling/broadcast construction when $d_{\mathrm{in}}\neq d_{\mathrm{out}}$ (and a learned scalar when $d_{\mathrm{in}}=d_{\mathrm{out}}$). The last columns count residual-path parameters only. For tasks with $d_{\mathrm{out}}>1$, MSE is computed as the mean of $(\hat{y}-y)^{2}$ over all output dimensions and samples.
@@ 3.2.1 Toy-$4\to 5$ synthetic regression task
Table 2: Ablation on Toy-2D-Vector. Mean test RMSE $\pm$ std over 10 seeds (5,000 epochs per run). Lower is better.
@@
Figure 5: The dual spline system in Sprecher Networks. Left: The inner spline $\phi$ is monotonic (non-decreasing) with fixed codomain $[0,1]$, parameterized via cumulative sums of softplus-transformed increments; it is strictly increasing on its spline domain and uses constant extension outside that domain (0 for inputs below the leftmost knot, 11 above the rightmost knot). Right: The outer spline $\Phi$ is a general (non-monotonic) univariate spline (piecewise-linear or cubic PCHIP) with optional learnable codomain parameters $(c_{c},c_{r})$ defining center and radius. It uses linear extrapolation outside its domain. Both domains can be updated during training as $\lambda$ and $\eta$ evolve; in some experiments we freeze domain updates after a warm-up period (and in our PINN experiments we disable domain updates altogether).
@@ 3.3 Normalization considerations
The unique structure of Sprecher Networks requires careful consideration when incorporating normalization techniques. While standard neural networks typically normalize individual neuron activations, the shared-spline architecture of SNs suggests a different approach: normalizing the entire vector output $\mathbf{h}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ of each Sprecher block as a cohesive unit.
@@
We often find it beneficial to skip normalization for the first block (default setting), allowing the network to directly process the input features while still stabilizing deeper layers. This selective application is particularly important when theoretical domain computation assumes inputs in $0,1]^{n}$, as normalization could shift the data outside expected ranges.
When computing theoretical domains (Section 7.4), the bounds must account for normalization:
  • Training mode: Conservative bounds assume standardization to approximately [4,4][-4,4] (covering  99.99% of a standard normal distribution), with affine transformation applied afterward.
  • Evaluation mode: Uses the actual per-channel running statistics (and affine parameters, if enabled) to compute tighter bounds.
@@ 3.4 Layer composition and final mapping
Let the input to the network be $\mathbf{h}^{(0)}=\mathbf{x}\in\mathbb{R}^{d_{0}}$ (where $d_{0}=d_{\mathrm{in}}$). The spline core of the $\ell$-th Sprecher block ($\ell=1,\dots,L$) produces a vector $\tilde{\mathbf{h}}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$, computed component-wise as:
$\tilde{\mathbf{h}}^{(\ell)}_{q}=\Phi^{(\ell)}\Biggl(\,\sum_{i=1}^{d_{\ell-1}}\lambda^{(\ell)}_{i}\,\phi^{(\ell)}\Bigl(\mathbf{h}^{(\ell-1)}_{i}+\eta^{(\ell)}\,q\Bigr)+\alpha q+\tau^{(\ell)}\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(\ell)}s_{j}^{(\ell)}\Biggr)+R^{(\ell)}\!\bigl(\mathbf{h}^{(\ell-1)}\bigr)_{q},\quad q=0,\dots,d_{\ell}-1.$ (2)
where the lateral mixing term is included when enabled, and $s_{j}^{(\ell)}$ denotes the pre-mixing activation for output $j$. The term $R^{(\ell)}(\mathbf{h}^{(\ell-1)})_{q}$ denotes an (optional) residual connection (e.g., the cyclic residual of Fig. 4 or a standard learned linear projection); set $R^{(\ell)}\equiv 0$ when residuals are disabled. When BatchNorm is used, it is applied before or after each block as described in Section 3.3. Throughout, $\tilde{\mathbf{h}}^{(\ell)}$ denotes the pre-normalization output of the Sprecher block; if BatchNorm-after-block is enabled then $\mathbf{h}^{(\ell)}$ denotes the post-normalization vector passed to the next layer (otherwise $\mathbf{h}^{(\ell)}=\tilde{\mathbf{h}}^{(\ell)}$).
@@
The network consists of exactly $L$ Sprecher blocks. The output of the final block, $\mathbf{h}^{(L)}\in\mathbb{R}^{d_{L}}$, is aggregated by summation to yield the scalar output:
@@
$\Bigl(T^{(\ell)}(z)\Bigr)_{q}=\Phi^{(\ell)}\Biggl(\,\sum_{i=1}^{d_{\ell-1}}\lambda^{(\ell)}_{i}\,\phi^{(\ell)}\Bigl(z_{i}+\eta^{(\ell)}\,q\Bigr)+\alpha q+\tau^{(\ell)}\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(\ell)}s_{j}^{(\ell)}\Biggr)+R^{(\ell)}(z)_{q},$
@@
This network uses $L$ blocks and $2L$ shared spline functions in total (one pair $(\phi^{(\ell)},\Phi^{(\ell)})$ per block).
@@
When the target function $f$ maps to $\mathbb{R}^{m}$ with $m>1$, the network first constructs the $L$ hidden layers as above, yielding a final hidden representation $\mathbf{h}^{(L)}\in\mathbb{R}^{d_{L}}$. An additional output block (block $L+1$) is then appended to map this representation $\mathbf{h}^{(L)}$ to the final output space $\mathbb{R}^{m}$. This $(L+1)$-th block operates without a final summation over its output index. It computes the final output vector $\mathbf{y}\in\mathbb{R}^{m}$ as:
@@
In summary: for an architecture with $L$ hidden layers, a scalar-output SN uses $L$ blocks and $2L$ shared splines. A vector-output SN (with $m>1$) uses $L+1$ blocks and $2(L+1)$ shared splines. This structure provides a natural extension of Sprecher’s original scalar formula to the vector-valued setting. In our supervised regression experiments we additionally apply a learned affine head to the final output,
$f_{\mathrm{final}}(\mathbf{x})\;=\;s_{\mathrm{out}}\,f(\mathbf{x})+b_{\mathrm{out}},$
with $s_{\mathrm{out}},b_{\mathrm{out}}\in\mathbb{R}$ (2 parameters, shared across output dimensions). We initialize $b_{\mathrm{out}}$ to the mean of the training targets (averaged over all outputs when $m>1$) and set $s_{\mathrm{out}}=0.1$. In PINN experiments (where no full supervised target set is available), we keep the default initialization $s_{\mathrm{out}}=1$ and $b_{\mathrm{out}}=0$.
@@
3.5 Illustrative expansions (scalar output)
To further clarify the compositional structure for the scalar output case ($m=1$), we write out the full expansions for networks with $L=1,2,3$ hidden layers, omitting optional residual connections and normalization operators for readability. For readability, we omit optional add-ons such as cyclic residual terms and normalization; when enabled, these appear additively (e.g., each block output includes an extra $R^{(\ell)}(\cdot)_{q}$ term as in (2)).
@@ 3.5.3 Three hidden layers ($L=3$)
$\mathbf{h}^{(2)}_{s}$ $=\Phi^{(2)}\Bigl(\sum_{r=0}^{d_{1}-1}\lambda^{(2)}_{r}\,\phi^{(2)}\Bigl(\mathbf{h}^{(1)}_{r}+\eta^{(2)}\,s\Bigr)+\alpha s+\tau^{(2)}\sum_{j\in\mathcal{N}(s)}\omega_{s,j}^{(2)}s_{j}^{(2)}\Bigr),\quad s=0,\dots,d_{2}-1,$
$\mathbf{h}^{(3)}_{q}$ $=\Phi^{(3)}\Bigl(\sum_{s=0}^{d_{2}-1}\lambda^{(3)}_{s}\,\phi^{(3)}\Bigl(\mathbf{h}^{(2)}_{s}+\eta^{(3)}\,q\Bigr)+\alpha q+\tau^{(3)}\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(3)}s_{j}^{(3)}\Bigr),\quad q=0,\dots,d_{3}-1.$
@@ 4 Comparison with related architectures
Parameter sharing None (typically) Activations? (Maybe) None (typically) Splines ($\phi,\Phi$) per block
@@
Theoretical basis UAT UAT KAT (inspired) KAS (Sprecher, direct)
@@
Memory complexity $O(LN^{2})$ $O(LN^{2})$ $O(LN^{2}G)$ $O(LN^{2})$ parallel
@@
Notes: LL=depth, NN=average width, GG=spline grid size/complexity. UAT=Universal Approx. Theorem, KAT=Kolmogorov-Arnold Theorem, KAS=Kolmogorov-Arnold-Sprecher. LAN details often follow KAN Appendix B [13]. Forward working memory refers to peak forward-pass intermediate storage plus parameters (excluding optimizer state), treating batch size as a constant; during training, reverse-mode autodiff may store additional activations unless checkpointing is used. The “sequential” mode for SNs processes output dimensions iteratively rather than in parallel, maintaining mathematical equivalence while reducing memory usage. All architectures may optionally include normalization layers; SNs apply normalization to entire block outputs rather than individual activations. For SNs, the $O(LN+LG)$ parameter scaling assumes the default cyclic/node residual; using a full linear residual projection adds an additional $O(\sum_{\ell}d_{\ell-1}d_{\ell})$ parameters. The parameter scaling notation uses $N$ to denote a typical or average layer width for simplicity, following [13]. For architectures with varying widths $d_{\ell}$, the $LN^{2}$ terms should be understood as $\sum_{\ell}d_{\ell-1}d_{\ell}$ (MLP, LAN), the $LN^{2}G$ term for KAN as $(\sum_{\ell}d_{\ell-1}d_{\ell})G$, and the $LN$ term for SN as $\sum_{\ell}(d_{\ell-1}+d_{\ell})$ (since SN uses weight vectors, not matrices), where the sum is over the relevant blocks/layers, for precise counts.
@@
The following proposition uses matrix weights $\lambda_{i,q}^{(\ell)}$ to establish the connection between SNs and LANs. However, the architecture we propose and analyze throughout this paper uses vector weights $\lambda_{i}^{(\ell)}$ (following Sprecher’s original formulation), where the same weights are shared across all output dimensions. This vector version represents an even more constrained special case of the LAN formulation shown below. The practical SN architecture with vector weights corresponds to the constraint that these matrix weights are constant across the output index, i.e., $\lambda_{i,q}^{(\ell)}=\lambda_{i}^{(\ell)}$ for all $q$. This constraint is fundamental to maintaining the ”true Sprecher” structure and achieving the characteristic $O(LN)$ parameter scaling. Additionally, when lateral mixing is enabled, the structure becomes more complex, requiring additional terms in the LAN representation to capture the intra-block communication.

### v2

@@ 3.2 Optional enhancements
  • Residual connections (node-wise / cyclic): When enabled, residual connections are implemented using a dimension-adaptive modulo-based cyclic assignment that maintains architectural coherence. We refer to this $O(\max(d_{\mathrm{in}},d_{\mathrm{out}}))$-parameter construction as a node-wise residual (in contrast to a dense linear residual projection). Like convolutional layers that achieve parameter efficiency through local connectivity patterns, cyclic residuals use periodic patterns to reduce parameters from $O(d_{\mathrm{in}}\times d_{\mathrm{out}})$ to $O(\max(d_{\mathrm{in}},d_{\mathrm{out}}))$. This method adapts based on dimensional relationships between adjacent layers:
@@
  • Output affine head: The network may include a learnable affine head applied to the final network output: $f_{\text{final}}(\mathbf{x})=s_{\text{out}}\,f(\mathbf{x})+b_{\text{out}}$, where $s_{\text{out}},b_{\text{out}}\in\mathbb{R}$ (2 parameters) are shared across all output dimensions. For supervised regression tasks we initialize $b_{\text{out}}$ to the mean of the training targets (averaged over all outputs when $m>1$) and set $s_{\text{out}}=0.1$; for classification and PINN experiments we keep the default initialization $s_{\text{out}}=1$ and $b_{\text{out}}=0$.
@@
Table 2: Cyclic residuals vs. linear residual projections on synthetic regression tasks. We report the best MSE on a fixed 1,024-point training set (measured in evaluation mode) achieved within the stated number of epochs. Toy-2D-Complex refers to the scalar 2D regression benchmark defined above; Toy-2D (vector) refers to the vector-valued 2D benchmark in Fig. 11; Toy-$4\to 5$ is a $0,1]^{4}\to\mathbb{R}^{5}$ regression task with outputs $f_{1}=\sin(2\pi x_{1})\cos(\pi x_{2})$, $f_{2}=\exp(-2(x_{1}^{2}+x_{2}^{2}))$, $f_{3}=x_{3}^{3}-x_{4}^{2}+\tfrac{1}{2}\sin(5x_{3})$, $f_{4}=\sigma(3(x_{1}+x_{2}-x_{3}+x_{4}))$, $f_{5}=\tfrac{1}{2}\sin(4\pi x_{1}x_{4})+\tfrac{1}{2}\cos(3\pi x_{2}x_{3})$ (where $\sigma$ denotes sigmoid). “Linear” uses a learned scalar when $d_{\mathrm{in}}=d_{\mathrm{out}}$ and a learned matrix $W\in\mathbb{R}^{d_{\mathrm{in}}\times d_{\mathrm{out}}}$ otherwise; “cyclic” uses the pooling/broadcast construction when $d_{\mathrm{in}}\neq d_{\mathrm{out}}$ (and a learned scalar when $d_{\mathrm{in}}=d_{\mathrm{out}}$). The last columns count residual-path parameters only. For tasks with $d_{\mathrm{out}}>1$, MSE is computed as the mean of $(\hat{y}-y)^{2}$ over all output dimensions and samples.
@@
Table 3: Ablation on the 2D vector-valued benchmark defined in Fig. 11 (“Toy-2D-Vector”). Mean test RMSE $\pm$ std over 10 seeds (5,000 epochs per run). Lower is better. Here, “+ domain tracking” updates spline domains without coefficient resampling, whereas “+ resampling” enables resampling.
@@
Figure 6: The dual spline system in Sprecher Networks. Left: The inner spline $\phi$ is monotonic (non-decreasing) with fixed codomain $[0,1]$, parameterized via cumulative sums of softplus-transformed increments; it is strictly increasing on its spline domain and uses constant extension outside that domain (0 for inputs below the leftmost knot, 11 above the rightmost knot). Right: The outer spline $\Phi$ is a general (non-monotonic) univariate spline (piecewise-linear or cubic PCHIP) with optional learnable codomain parameters $(c_{c},c_{r})$ defining center and radius. It uses linear extrapolation outside its domain. Both domains can be updated during training as $\lambda$ and $\eta$ evolve; unless explicitly stated otherwise, we perform dynamic domain updates during training (the PINN setting in Sec. 8 is an exception: spline domains are kept fixed (no domain updates), as described there).
@@ 3.3 Normalization considerations
The unique structure of Sprecher Networks requires careful consideration when incorporating normalization techniques. While standard neural networks typically normalize individual neuron activations, the shared-spline architecture of SNs suggests applying normalization only at block boundaries: we normalize the block output vector $\mathbf{h}^{(\ell)}\in\mathbb{R}^{d_{\ell}}$ (featurewise, as in standard BatchNorm) rather than normalizing the internal scalar pre-activations within a block.
@@
We often find it beneficial to skip normalization for the first block (default setting), allowing the network to directly process the input features while still stabilizing deeper layers. This also preserves the canonical assumption that raw inputs lie in $0,1]^{n}$; if normalization is applied before the first block, bound propagation simply starts from the correspondingly transformed input interval.
When computing theoretical domains (Section 7.5), the bounds must account for normalization, since subsequent $\phi^{(\ell)}$ splines receive post-normalization activations:
  • Batch-statistics mode (training, and also the “batch-stat” evaluation protocol used in several of our experiments): Conservative bounds assume standardization to approximately [4,4][-4,4] (covering  99.99% of a standard normal distribution); this is a heuristic bound, with affine transformation applied afterward.
  • Running-statistics mode (standard BatchNorm evaluation): Uses the stored per-channel running statistics (and affine parameters, if enabled) to compute tighter bounds.
This means spline-domain updates track the post-normalization ranges, so normalization is explicitly accounted for when setting subsequent spline domains.
In theoretical domain propagation, the interval passed to the next block is always taken after applying the normalization map (whether normalization is placed before or after a block), so subsequent $\phi$ splines do not “miss” values due to normalization.
@@ 3.4 Layer composition and final mapping
The $\ell$-th block map is given in Eq. (2) (introduced above). Throughout, we write $\tilde{\mathbf{h}}^{(\ell)}$ for the pre-normalization output of block $\ell$ (after $\Phi^{(\ell)}$ and any residual term), and $\mathbf{h}^{(\ell)}$ for the vector passed to the next block after any optional normalization (otherwise $\mathbf{h}^{(\ell)}=\tilde{\mathbf{h}}^{(\ell)}$). More explicitly, the (unmixed) pre-activation for output $q$ is
$s_{q}^{(\ell)}:=\sum_{i=1}^{d_{\ell-1}}\lambda^{(\ell)}_{i}\,\phi^{(\ell)}\Bigl(\mathbf{h}^{(\ell-1)}_{i}+\eta^{(\ell)}\,q\Bigr)+\alpha q,$
and the lateral mixing adds a weighted combination of neighboring $s_{j}^{(\ell)}$ values before applying $\Phi^{(\ell)}$. The term $R^{(\ell)}(\mathbf{h}^{(\ell-1)})_{q}$ denotes an (optional) residual connection (e.g., the node-wise/cyclic residual of Fig. 5 or a standard learned linear projection); set $R^{(\ell)}\equiv 0$ when residuals are disabled. Recall that $\tilde{\mathbf{h}}^{(\ell)}$ denotes the pre-normalization block output and $\mathbf{h}^{(\ell)}$ the post-normalization vector (Section 3.3).
@@
One can construct a scalar-output SN using exactly $L$ Sprecher blocks. The output of the final block, $\mathbf{h}^{(L)}\in\mathbb{R}^{d_{L}}$, is aggregated by summation to yield the scalar output:
@@
$\Bigl(T^{(\ell)}(\mathbf{z})\Bigr)_{q}=\Phi^{(\ell)}\Biggl(\,\sum_{i}\lambda^{(\ell)}_{i}\,\phi^{(\ell)}\Bigl(z_{i}+\eta^{(\ell)}\,q\Bigr)+\alpha q+\tau^{(\ell)}\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(\ell)}s_{j}^{(\ell)}\Biggr)+R^{(\ell)}(\mathbf{z})_{q},$ (3)
@@
(The summation index $i$ ranges over the input coordinates to block $\ell$: for the first block, $i=1,\ldots,d_{0}$ following Sprecher’s 1-based convention for raw inputs; for subsequent blocks, $i$ ranges over the preceding block’s outputs using 0-based indexing as noted in Section 2.) This summed scalar-output form uses $L$ blocks and $2L$ shared spline functions in total (one pair $(\phi^{(\ell)},\Phi^{(\ell)})$ per block). Unless explicitly stated otherwise, our scalar-output experiments use this summed form. In Section 1.1 we instead use the $m=1$ vector-valued construction below, appending an additional non-summed output block with $d^{(L+1)}=1$, to align depth counting with the MLP baselines.
@@
When the target function $f$ maps to $\mathbb{R}^{m}$ with $m>1$, the network first constructs the $L$ hidden layers as above, yielding a final hidden representation $\mathbf{h}^{(L)}\in\mathbb{R}^{d_{L}}$. A final output block (block $L+1$) is then appended to map this representation $\mathbf{h}^{(L)}$ to the final output space $\mathbb{R}^{m}$. This $(L+1)$-th block operates without a final summation over its output index. It computes the final output vector $\mathbf{y}\in\mathbb{R}^{m}$ as:
@@
Summary. For an architecture with $L$ hidden layers, the sum-aggregated scalar-output SN uses $L$ blocks and $2L$ shared splines. A vector-output SN (with $m>1$) uses $L+1$ blocks and $2(L+1)$ shared splines. This structure provides a natural extension of Sprecher’s original scalar formula to the vector-valued setting. In our experiments we apply the learned affine head described in Section 3.2 to the final output.
@@
3.5 Illustrative expansions (sum-aggregated scalar output)
To further clarify the compositional structure for the sum-aggregated scalar-output case ($m=1$), we write out the full expansions for networks with $L=1,2,3$ hidden layers. For conciseness, we omit optional add-ons such as cyclic residual terms and normalization; when enabled, these appear additively (e.g., each block output includes an extra $R^{(\ell)}(\cdot)_{q}$ term as in (3)).
@@ 3.5.3 Three hidden layers ($L=3$)
$\mathbf{h}^{(2)}_{p}$ $=\Phi^{(2)}\Bigl(\sum_{r=0}^{d_{1}-1}\lambda^{(2)}_{r}\,\phi^{(2)}\Bigl(\mathbf{h}^{(1)}_{r}+\eta^{(2)}\,p\Bigr)+\alpha p+\tau^{(2)}\sum_{j\in\mathcal{N}(p)}\omega_{p,j}^{(2)}s_{j}^{(2)}\Bigr),\quad p=0,\dots,d_{2}-1,$
$\mathbf{h}^{(3)}_{q}$ $=\Phi^{(3)}\Bigl(\sum_{p=0}^{d_{2}-1}\lambda^{(3)}_{p}\,\phi^{(3)}\Bigl(\mathbf{h}^{(2)}_{p}+\eta^{(3)}\,q\Bigr)+\alpha q+\tau^{(3)}\sum_{j\in\mathcal{N}(q)}\omega_{q,j}^{(3)}s_{j}^{(3)}\Bigr),\quad q=0,\dots,d_{3}-1.$
@@ 4 Comparison with related architectures
Parameter sharing None (typically) Activations (optional) None (typically) Splines ($\phi,\Phi$) per block
@@
Theoretical basis UAT UAT KAT (inspired) Sprecher (direct)
@@
Memory complexity $O(LN^{2})$ $O(LN^{2}+LNG)$ $O(LN^{2}G)$ $O(LN^{2})$ parallel
@@
Notes: LL=depth, NN=average width, GG=spline grid size/complexity. UAT=Universal Approx. Theorem, KAT=Kolmogorov-Arnold Theorem. LAN details often follow KAN Appendix B [12]. Forward working memory refers to peak forward-pass intermediate storage plus parameters (excluding optimizer state), treating batch size as a constant; during training, reverse-mode autodiff may store additional activations unless checkpointing is used. The “sequential” mode for SNs processes output dimensions iteratively rather than in parallel, maintaining mathematical equivalence while reducing memory usage. All architectures may optionally include normalization layers; SNs apply normalization to entire block outputs rather than individual activations. For SNs, the $O(LN+LG)$ parameter scaling assumes the cyclic/node residual; using a full linear residual projection adds an additional $O(\sum_{\ell}d_{\ell-1}d_{\ell})$ parameters. The parameter scaling notation uses $N$ to denote a typical or average layer width for simplicity, following [12]. For architectures with varying widths $d_{\ell}$, the $LN^{2}$ terms should be understood as $\sum_{\ell}d_{\ell-1}d_{\ell}$ (MLP, LAN), the $LN^{2}G$ term for KAN as $(\sum_{\ell}d_{\ell-1}d_{\ell})G$, and the $LN$ term for SN as $\sum_{\ell}(d_{\ell-1}+d_{\ell})$ (since SN uses weight vectors, not matrices), where the sum is over the relevant blocks/layers, for precise counts.
@@
The following proposition uses matrix weights $\lambda_{i,q}^{(\ell)}$ to establish the connection between SNs and LANs. However, the architecture we propose and analyze throughout this paper uses vector weights $\lambda_{i}^{(\ell)}$ (following Sprecher’s original formulation), where the same weights are shared across all output dimensions. This vector version represents an even more constrained special case of the LAN formulation shown below. The practical SN architecture with vector weights corresponds to the constraint that these matrix weights are constant across the output index, i.e., $\lambda_{i,q}^{(\ell)}=\lambda_{i}^{(\ell)}$ for all $q$. This constraint is fundamental to maintaining the “true Sprecher” structure and achieving the characteristic $O(LN)$ parameter scaling. Additionally, when lateral mixing is enabled, the structure becomes more complex, requiring additional terms in the LAN representation to capture the intra-block communication.